What is the sum of all odd integers between 8 and 26?
Answer by ikleyn(52847)

.

There are no so many such odd integers.


You can write them in one line  9 11  13  15  17  19  21  23  25


and add them, by grouping


    9 + 11 = 20

   13 + 17 = 30

   15 + 25 = 40

   19 + 21 = 40


So the total sum is  20 + 30 + 40 + 40 + 23 = 153.      ANSWER

Solved.

Good problem for an advanced kindergarten alumni.